# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@
Abstract.
We revisit an identification of the quantum Toda lattice for $\mathrm{GL}_{N}$ and the truncated shifted Yangian of $\mathfrak{sl}_{2}$, as well as related constructions, from a purely algebraic point of view, bypassing the topological medium of the homology of the affine Grassmannian. For instance, we interpret the Gerasimov-Kharchev-Lebedev-Oblezin homomorphism into the algebra of difference operators via a finite analog of the Miura transform. This algebraic identification is deduced by relating degenerate affine Hecke algebras to the simplest example of a rational Feigin-Odesskii shuffle product. As a bonus, we obtain a presentation of the latter via a mirabolic version of the Kostant-Whittaker reduction.
Introduction
Shuffle algebras are ubiquitous in representation theory. To name a few applications, they are related to infinite-dimensional quantum groups (including quantum toroidal algebras), Hall algebras, canonical bases, and so on. They were introduced by Feigin-Odesskii in a series of papers [OF89, OF97, FO98]. In the simplest case, a shuffle algebra is, as a vector space, the direct sum $\bigoplus_{k}\mathcal{O}(Y_{1},\ldots,Y_{k})^{S_{k}}$ of suitably defined symmetric functions in $k$ variables, with the product
$\mathcal{O}(Y_{1},\ldots,Y_{k})^{S_{k}}$ $\times\mathcal{O}(Y_{1},\ldots,Y_{l})^{S_{l}}\rightarrow\mathcal{O}(Y_{1},\ldots,Y_{k+l})^{S_{k+l}},$
$(f*g)(Y_{1},\ldots,Y_{k+l})$ $=\frac{1}{k!l!}\mathrm{Sym}_{k+l}\left(\prod_{i<j}\frac{\theta(Y_{i}-Y_{j}+1)}{\theta(Y_{i}-Y_{j})}f(Y_{1},\ldots,Y_{k})g(Y_{k+1},\ldots,Y_{k+l})\right).$
@@ Introduction
(1) $\theta^{\mathrm{ell}}(u)=\vartheta_{1}(u|\tau),\qquad\theta^{\mathrm{trig}}(u)=\sin(\pi u/\tau),\qquad\theta^{\mathrm{rat}}(u)=u,$
@@
$\mathcal{O}(Y_{1},\ldots,Y_{k})^{S_{k}}={}^{N}\mathcal{S}^{\mathrm{ell}}_{k}:=S^{k}H^{0}(\Theta),\qquad{}^{N}\mathcal{S}^{\mathrm{ell}}=\bigoplus_{k}{}^{N}\mathcal{S}^{\mathrm{ell}}_{k},$
@@
In this paper, we study the rational case. The elliptic one above admits a degeneration to it [EH04] so that
(2) $^{N}\mathcal{S}^{\mathrm{rat}}=\bigoplus_{k}{}^{N}\mathcal{S}_{k}^{\mathrm{rat}},\qquad{}^{N}\mathcal{S}^{\mathrm{rat}}_{k}=\mathbf{C}[Y_{1},\ldots,Y_{k}]^{S_{k}}_{<N},$
where $\mathbf{C}[Y_{1},\ldots,Y_{k}]^{S_{k}}_{<N}$ is the space of symmetric polynomials of degree less than $N$ in each variable, and the factor is $\theta^{\mathrm{rat}}(u)$. More generally, one can drop the restriction on the degree and consider $\mathcal{S}^{\mathrm{rat}}=\bigoplus_{k}\mathcal{S}^{\mathrm{rat}}_{k}$ with $\mathcal{S}^{\mathrm{rat}}_{k}=\mathbf{C}[Y_{1},\ldots,Y_{k}]^{S_{k}}$.
The first goal of the paper is to relate these shuffle algebras to degenerate affine Hecke algebras whose definition we recall in Section˜2.2. The idea is as follows: observe that without the theta factor, the shuffle product is simply the multiplication on the symmetric algebra of the vector space $\mathbf{C}[Y_{1}]$ or its $N$-th truncation. Instead of deforming it via inserting a factor, we can deform the symmetrization itself, that is, the action of the symmetric group. One such deformation is provided by the Demazure operators, so that a simple transposition acts on $\mathbf{C}[Y_{1},\ldots,Y_{k}]$ by
@@
where $P_{i,i+1}$ permutes variables $Y_{i}$ and $Y_{i+1}$. Denoting by $e_{k}\in\mathbf{C}[S_{k}]$ the symmetrizer idempotent in the group algebra of the symmetric group, this is the action of $S_{k}$ on the spherical module $\mathfrak{H}_{k}e_{k}\cong\mathbf{C}[Y_{1},\ldots,Y_{k}]$ of the degenerate affine Hecke algebra $\mathfrak{H}_{k}$. Its spherical subalgebra $e_{k}\mathfrak{H}_{k}e_{k}$ can be identified with the symmetric polynomials $\mathcal{S}^{\mathrm{rat}}_{k}$.
@@
$(f,g)\mapsto f*^{\mathrm{Hecke}}g:=e_{k+l}f(Y_{1},\ldots,Y_{k})g(Y_{k+1},\ldots,Y_{k+l})e_{k+l}\in e_{k+l}\mathfrak{H}_{k+l}e_{k+l}\cong\mathcal{S}^{\mathrm{rat}}_{k+l}.$
The main result of the first part of the paper is Theorem˜2.11.
@@
The rational Feigin-Odesskii shuffle product coincides with the one induced from degenerate affine Hecke algebras.
We must admit that the result is by no means new: for instance, its trigonometric version can considered as simply a reformulation of the known relation between explicit formulas for Hall-Littlewood polynomials and affine Hecke algebras [Mac95]. Our main contribution is an alternative proof via the Shibukawa-Ueno $R$-operators whose definition we recall in Section˜2.1. Those are particular solutions of the quantum Yang-Baxter equation
@@
introduced in [SU92] that act on the space of suitably defined functions and take as an input any solution to the three-term relation (2.1) (in particular, one of the functions (1)). At the same time, there is a canonical affinization procedure [KS97] that produces a solution to the quantum Yang-Baxter equation out of a representation of a symmetric group (or a finite Hecke algebra). It is an easy calculation that the affinization of the Demazure action above is essentially the rational Shibukawa-Ueno $R$-operator (and, similarly, the affinization of the action of the finite Hecke algebra on the spherical module of the affine Hecke algebra is its trigonometric version, although we do not calculate it in the paper and leave it as remark). Then the explicit formula for the symmetrization of the Demazure action can be computed via a certain combination of the $R$-operators which is done in Proposition˜2.3.
@@
Interestingly, as we mentioned, there is an elliptic version of the Shibukawa-Ueno operator. Unfortunately, we do not know if and how it might be related to elliptic affine Hecke algebras [GKV97]. Curiously, the elliptic case admits a dual description on the $\mathrm{GL}_{N}$-side, in the sense of Fourier-Mukai transform on an elliptic curve. Namely, by [FP94], the elliptic Shibukawa-Ueno $R$-operator is gauge-equivalent to Belavin’s $R$-matrix [Bel80, Che86] whose quasi-periodic properties can be interpreted via an essentially unique stable vector bundle of rank $N$ and degree 1 on an elliptic curve. A natural question is: what is the meaning of the shuffle product on the $\mathrm{GL}_{N}$-side?
@@
We approach this question for the rational case in the second part of the paper where we apply the construction of the first part to the quantum Toda lattice for $\mathrm{GL}_{N}$. The idea to use degenerate affine Hecke algebras is inspired by the results of [CK18] where the authors constructed and proved a K-theoretic version of the geometric Satake equivalence for a certain subcategory of the $q$-analog of Harish-Chandra bimodules (we recall the rational versions below). Following the Kazhdan-Kostant formulation of this integrable system [Kos78], let us consider a non-degenerate character $\psi$ of the Lie algebra $\mathfrak{n}_{+}$ of the subgroup of upper unitriangular matrices $N_{+}$. The action of $\mathfrak{n}_{+}$ via right- and left-invariant vector fields gives rise to a homomorphism $\mathrm{U}(\mathfrak{n}_{+})\otimes\mathrm{U}(\mathfrak{n}_{+})\rightarrow\mathrm{D}(\mathrm{GL}_{N})$ into the algebra of differential operators. Then the Toda Hamiltonians are realized as the image of bi-invariant differential operators, which can be identified with the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}(\mathfrak{gl}_{N})$, inside the quantum Hamiltonian reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ with respect to the latter embedding and the character $(-\psi,\psi)$. This is an example of the Kostant-Whittaker reduction for Harish-Chandra bimodules studied in [BF08].
@@
According to [BFN19, Appendix B], the Toda lattice can be identified with a certain quotient $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})$ of the shifted Yangian $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ whose definition we recall in Section˜2.3. For introduction, we will only use the fact that it is generated by the coefficients of the series
@@
and that the coefficients of $x_{+}(u)$ generate the subalgebra isomorphic to the shuffle algebra $\mathcal{S}^{\mathrm{rat}}$ [Tsy21]. Therefore, it gives a $\mathrm{GL}_{N}$-presentation of $\mathcal{S}^{\mathrm{rat}}$. However, citing [Fin+18], “we must admit that the identification of the quantum Toda [lattice] for $\mathrm{GL}_{N}$ and the [truncated] shifted Yangian for $\mathfrak{sl}_{2}$ (purely algebraic objects) goes through a topological medium: equivariant homology of the affine Grassmannian of $\mathrm{GL}_{N}$”; this additional step is the derived Satake equivalence of [BF08]. For instance, while $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ is defined via generators and relations, this identification does not specify explicitly the images of the former.
@@
as well as related constructions, in a purely algebraic and explicit way, bypassing the topological medium, and “refine” it to a mirabolic version of the Kostant-Whittaker reduction. It is based on another classical fact. For any $\mathrm{GL}_{N}$-representation $V$, the free left $\mathrm{U}(\mathfrak{gl}_{N})$-module $\mathrm{U}(\mathfrak{gl}_{N})\otimes V$ has a right $\mathrm{U}(\mathfrak{gl}_{N})$ action defined from the diagonal group action of $\mathrm{GL}_{N}$ on the tensor product. This is an example of a (free) Harish-Chandra bimodule whose definition and properties we review in Section˜3. For the tensor product of the vector representation $V=(\mathbf{C}^{N})^{\otimes k}$, there is an action of the degenerate affine Hecke algebra $\mathfrak{H}_{k}$ on $\mathrm{U}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ respecting the $\mathrm{U}(\mathfrak{gl}_{N})$-bimodule structure. In particular, its spherical subalgebra $\mathcal{S}_{k}\cong e_{k}\mathfrak{H}_{k}e_{k}$ acts on the $S_{k}$-invariants that can be identified with the free Harish-Chandra bimodule $\mathrm{U}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ associated to the $k$-th symmetric power. By restricting the action on the highest-weight vector for all $k$, one obtains a map of vector spaces
@@
where the target is the quantum Hamiltonian reduction of the semi-direct product with $\mathfrak{gl}_{N}$ acting on the symmetric algebra $S^{\bullet}\mathbf{C}^{N}$ in the standard way. In particular, it is an algebra as well. If we denote by $\mathfrak{m}_{N}$ the mirabolic subalgebra of the subgroup of $\mathrm{GL}_{N}$ preserving the dual of the last basis vector in $\mathbf{C}^{N}$ (see Section˜5.3 for precise definition), then it turns out that this map restricts to a similarly defined quantum Hamiltonian reduction
(3) $^{N}\mathcal{S}^{\mathrm{rat}}\rightarrow\mathrm{U}(\mathfrak{m}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$
The following is Theorem˜5.14 and Theorem˜5.9; here $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})^{\geq 0}$ is the non-negative part of the shifted Yangian. In fact, the explicit formulas for the second homomorphism are close to another presentation of the quantization of Zastava spaces for $\mathfrak{sl}_{2}$ from [FR14, Section 3.22].
@@
The map $^{N}\mathcal{S}^{\mathrm{rat}}\xrightarrow{\sim}\mathrm{U}(\mathfrak{m}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ is an algebra isomorphism. It can be extended to a surjective algebra homomorphism $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})^{\geq 0}\rightarrow\mathrm{U}(\mathfrak{gl}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$.
Unfortunately, we do not know if and how this result can be related to the aforementioned stable bundles on an elliptic curve. We expect (or, rather, hope) that it could viewed as a quantum rational version of a certain algebra of Hecke modifications of semistable bundles of degree 0, see Remark˜5.15. In general, it would be interesting to interpret this result at least on the classical level in terms of elliptic analog of the Kostant slice [FM00].
Besides degenerate affine Hecke algebras, the other main tool to prove the theorem is the Kirillov projector $P_{\psi}$ introduced in [Kal25]. This is an element in a certain completion of $\mathrm{U}(\mathfrak{m}_{N})$ that acts on Whittaker modules and projects onto the space of Whittaker vectors (that is, on which $\mathfrak{n}_{+}$ acts with character $\psi$); as such, it is an analog of the classical extremal projector [AST71] in the case of trivial character. For instance, the Kirillov projector gives a canonical isomorphism $V\xrightarrow{\sim}\mathrm{U}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+}$ of any $\mathrm{GL}_{N}$-representation $V$ with the mirabolic version of the Kostant-Whittaker reduction. It turns out that this isomorphism has rather peculiar properties with respect to the degenerate affine Hecke algebra action mentioned above, see Section˜5.3. For example, under the isomorphism (3), the standard monomial basis of $S^{\bullet}\mathbf{C}^{N}$ is mapped to a rational version of the Hall-Littlewood polynomials, see Example˜5.12.
We extend this homomorphism to the Toda lattice. Naturally, $\mathrm{GL}_{N}$ is embedded into the matrix space; denote by $X=(x_{ji})$ the (transposed) matrix with $x_{ij}\in\mathcal{O}(\mathrm{GL}_{N})$ the $(i,j)$-matrix entry function. Its inverse $X^{-1}$ is well-defined in $\mathcal{O}(\mathrm{GL}_{N})$. Also, denote by $E^{R}=(E_{ij}^{R})$ the matrix composed of left-invariant vector fields associated to the standard generators of the Lie algebra $\mathfrak{gl}_{N}$. The following is Theorem˜7.9. It would interesting to compare this explicit form with the approach of [Fin+18] via coproducts on both sides.
@@
$x_{+}(u)$ $\mapsto-\sum_{i=0}^{\infty}[(E^{R})^{i}X]_{1N}u^{-i-1},\qquad x_{-}(u)\mapsto\sum_{i=0}^{\infty}[X^{-1}(E^{R})^{i}]_{1N}u^{-i-1},$
@@
The following is Theorem˜7.7 and the second part of Theorem˜7.9

### v4

@@
[…]
Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice
[…]
Shuffle Products, Degenerate Affine Hecke Algebras,
and Quantum Toda Lattice
[…]
Artem KALMYKOV
[…]
A. Kalmykov
[…]
Department of Mathematics and Statistics, McGill University,
[…]
Received August 14, 2025, in final form February 09, 2026; Published online February 23, 2026
[…]
We revisit an identification of the quantum Toda lattice for $\mathrm{GL}_{N}$ and the truncated shifted Yangian of $\mathfrak{sl}_{2}$, as well as related constructions, from a purely algebraic point of view, bypassing the topological medium of the homology of the affine Grassmannian. For instance, we interpret the Gerasimov–Kharchev–Lebedev–Oblezin homomorphism into the algebra of difference operators via a finite analog of the Miura transform. This algebraic identification is deduced by relating degenerate affine Hecke algebras to the simplest example of a rational Feigin–Odesskii shuffle product. As a bonus, we obtain a presentation of the latter via a mirabolic version of the Kostant–Whittaker reduction.
[…]
Yangian; Toda lattice; shuffle algebra; affine Hecke algebra
[…]
17B37; 17B38
1 Introduction
Shuffle algebras are ubiquitous in representation theory. To name a few applications, they are related to infinite-dimensional quantum groups (including quantum toroidal algebras), Hall algebras, canonical bases, and so on. They were introduced by Feigin–Odesskii in a series of papers [12, 40, 41]. In the simplest case, a shuffle algebra is, as a vector space, the direct sum $\bigoplus_{k}\mathcal{O}(Y_{1},\dots,Y_{k})^{S_{k}}$ of suitably defined symmetric functions in $k$ variables, with the product
$\mathcal{O}(Y_{1},\dots,Y_{k})^{S_{k}}\times\mathcal{O}(Y_{1},\dots,Y_{l})^{S_{l}}\rightarrow\mathcal{O}(Y_{1},\dots,Y_{k+l})^{S_{k+l}},$ (1.1)
$(f*g)(Y_{1},\dots,Y_{k+l})=\frac{1}{(k+l)!}\mathrm{Sym}_{k+l}\biggl(\prod_{i<j}\!\frac{\theta(Y_{i}-Y_{j}+1)}{\theta(Y_{i}\!-\!Y_{j})}f(Y_{1},\dots,Y_{k})g(Y_{k+1},\dots,Y_{k+l})\!\biggr),$
@@ 1 Introduction
$\theta^{\mathrm{ell}}(u)=\vartheta_{1}(u|\tau),\qquad\theta^{\mathrm{trig}}(u)=\sin(\pi u/\tau),\qquad\theta^{\mathrm{rat}}(u)=u,$ (1.2)
@@
$\mathcal{O}(Y_{1},\dots,Y_{k})^{S_{k}}={}^{N}\!\mathcal{S}^{\mathrm{ell}}_{k}:=S^{k}H^{0}(\Theta),\qquad{}^{N}\!\mathcal{S}^{\mathrm{ell}}=\bigoplus_{k}{}^{N}\!\mathcal{S}^{\mathrm{ell}}_{k},$
@@
In this paper, we study the rational case. The elliptic one above admits a degeneration to it [10] so that
$^{N}\!\mathcal{S}^{\mathrm{rat}}=\bigoplus_{k}{}^{N}\!\mathcal{S}_{k}^{\mathrm{rat}},\qquad{}^{N}\!\mathcal{S}^{\mathrm{rat}}_{k}=\mathbf{C}[Y_{1},\dots,Y_{k}]^{S_{k}}_{<N},$
where $\mathbf{C}[Y_{1},\dots,Y_{k}]^{S_{k}}_{<N}$ is the space of symmetric polynomials of degree less than $N$ in each variable, and the factor is $\theta^{\mathrm{rat}}(u)$. More generally, one can drop the restriction on the degree and consider $\mathcal{S}^{\mathrm{rat}}=\bigoplus_{k}\mathcal{S}^{\mathrm{rat}}_{k}$ with $\mathcal{S}^{\mathrm{rat}}_{k}=\mathbf{C}[Y_{1},\dots,Y_{k}]^{S_{k}}$.
The first goal of the paper is to relate these shuffle algebras to degenerate affine Hecke algebras whose definition we recall in Section 3.2. The idea is as follows: observe that without the theta factor, the shuffle product is simply the multiplication on the symmetric algebra of the vector space $\mathbf{C}[Y_{1}]$ or its $N$-th truncation. Instead of deforming it via inserting a factor, we can deform the symmetrization itself, that is, the action of the symmetric group. One such deformation is provided by the Demazure operators, so that a simple transposition acts on $\mathbf{C}[Y_{1},\dots,Y_{k}]$ by
@@
where $P_{i,i+1}$ permutes variables $Y_{i}$ and $Y_{i+1}$. Denoting by $e_{k}\in\mathbf{C}[S_{k}]$ the symmetrizer idempotent in the group algebra of the symmetric group, this is the action of $S_{k}$ on the spherical module $\mathfrak{H}_{k}e_{k}\cong\mathbf{C}[Y_{1},\dots,Y_{k}]$ of the degenerate affine Hecke algebra $\mathfrak{H}_{k}$. Its spherical subalgebra $e_{k}\mathfrak{H}_{k}e_{k}$ can be identified with the symmetric polynomials $\mathcal{S}^{\mathrm{rat}}_{k}$.
@@
$(f,g)\mapsto f*^{\mathrm{Hecke}}g:=e_{k+l}f(Y_{1},\dots,Y_{k})g(Y_{k+1},\dots,Y_{k+l})e_{k+l}\in e_{k+l}\mathfrak{H}_{k+l}e_{k+l}\cong\mathcal{S}^{\mathrm{rat}}_{k+l}.$
The following result of Theorem 3.11 seems to be folklore, but, unfortunately, we were not able to find a reference for its algebraic proof. A geometric proof can be found in an unpublished preprint by Grojnowski [23].
@@
The rational Feigin–Odesskii shuffle product coincides with the one induced from degenerate affine Hecke algebras.
In principle, the theorem follows directly from the explicit formula for the symmetrizer in the spherical representation of a degenerate affine Hecke algebra and, strictly speaking, is not new: for instance, in the trigonometric case, it reflects a well-known relation between Hall–Littlewood polynomials and affine Hecke algebras [36]. Our main contribution is an alternative proof based on the following simple observation: a canonical affinization procedure [32], which, in general, produces a solution of a quantum Yang–Baxter equation
@@
out of a representation of symmetric group (or, in the trigonometric case, of a finite Hecke algebra), applied to the spherical module gives a particular example of a Shibukawa–Ueno $R$-operator introduced in [44] whose definition we recall in Section 3.1. In general, the latter acts on the space of suitably defined functions and takes as an input any solution to the three-term relation (3.1); in particular, the case of the spherical module of a degenerate affine Hecke algebra corresponds to the rational factor in (1.2), while that of a non-degenerate affine Hecke algebra – to the trigonometric one. Then the theorem is deduced by, on the one hand, representing the symmetrizer though a certain combination of $R$-matrices, on the other hand, expressing this combination as the usual symmetrization map, but with a factor akin to (1.1), see Proposition 3.3.
The main advantage of this proof is that the second part works uniformly for any factor that satisfies the three-term relation. In principle, there are (formal) analogs of theta functions for arbitrary generalized cohomology theory with orientation [42]. Unfortunately, we do not know how the respective Shibukawa–Ueno $R$-operators are related to the corresponding formal (affine) Hecke algebras [24], however, we believe that similar formulas should hold in more general cases.
@@
In particular, it would be interesting to study the elliptic version of the Shibukawa–Ueno operator and its potential relation to elliptic affine Hecke algebras [21]. Curiously, the elliptic case admits a dual description on the $\mathrm{GL}_{N}$-side, in the sense of Fourier–Mukai transform on an elliptic curve. Namely, by [13], the elliptic Shibukawa–Ueno $R$-operator is gauge-equivalent to Belavin’s $R$-matrix [2, 8] whose quasi-periodic properties can be interpreted via an essentially unique stable vector bundle of rank $N$ and degree 1 on an elliptic curve. A natural question is: what is the meaning of the shuffle product on the $\mathrm{GL}_{N}$-side?
@@
We approach this question for the rational case in the second part of the paper where we apply the construction of the first part to the quantum Toda lattice for $\mathrm{GL}_{N}$. The idea to use degenerate affine Hecke algebras is inspired by the results of [5] where the authors constructed and proved a K-theoretic version of the geometric Satake equivalence for a certain subcategory of the $q$-analog of Harish-Chandra bimodules (we recall the rational versions below). Following the Kazhdan–Kostant formulation of this integrable system [33], let us consider a non-degenerate character $\psi$ of the Lie algebra $\mathfrak{n}_{+}$ of the subgroup of upper unitriangular matrices $N_{+}$. The action of $\mathfrak{n}_{+}$ via right- and left-invariant vector fields gives rise to a homomorphism $\mathrm{U}(\mathfrak{n}_{+})\otimes\mathrm{U}(\mathfrak{n}_{+})\rightarrow\mathrm{D}(\mathrm{GL}_{N})$} into the algebra of differential operators. Then the Toda Hamiltonians are realized as the image of bi-invariant differential operators, which can be identified with the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}(\mathfrak{gl}_{N})$, inside the quantum Hamiltonian reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ with respect to the latter embedding and the character $(-\psi,\psi)$. This is an example of the Kostant–Whittaker reduction for Harish-Chandra bimodules studied in [3].
@@
According to [4, Appendix B], the Toda lattice can be identified with a certain quotient $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})$ of the shifted Yangian $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ whose definition we recall in Section 3.3. For introduction, we will only use the fact that it is generated by the coefficients of the series
@@
and that the coefficients of $x_{+}(u)$ generate the subalgebra isomorphic to the shuffle algebra $\mathcal{S}^{\mathrm{rat}}$ [45]. Therefore, it gives a $\mathrm{GL}_{N}$-presentation of $\mathcal{S}^{\mathrm{rat}}$. However, citing [14], “we must admit that the identification of the quantum Toda [lattice] for $\mathrm{GL}_{N}$ and the [truncated] shifted Yangian for $\mathfrak{sl}_{2}$ (purely algebraic objects) goes through a topological medium: equivariant homology of the affine Grassmannian of $\mathrm{GL}_{N}$”; this additional step is the derived Satake equivalence of [3]. For instance, while $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})$ is defined via generators and relations, this identification does not specify explicitly the images of the former.
@@
as well as related constructions, in a purely algebraic and explicit way, bypassing the topological medium, and “refine” it to a mirabolic version of the Kostant–Whittaker reduction. It is based on another classical fact. For any $\mathrm{GL}_{N}$-representation $V$, the free left $\mathrm{U}(\mathfrak{gl}_{N})$-module $\mathrm{U}(\mathfrak{gl}_{N})\otimes V$ has a right $\mathrm{U}(\mathfrak{gl}_{N})$ action defined from the diagonal group action of $\mathrm{GL}_{N}$ on the tensor product. This is an example of a (free) Harish-Chandra bimodule whose definition and properties we review in Section 4. For the tensor product of the vector representation $V=\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$, there is an action of the degenerate affine Hecke algebra $\mathfrak{H}_{k}$ on $\mathrm{U}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$ respecting the $\mathrm{U}(\mathfrak{gl}_{N})$-bimodule structure. In particular, its spherical subalgebra $\mathcal{S}_{k}\cong e_{k}\mathfrak{H}_{k}e_{k}$ acts on the $S_{k}$-invariants that can be identified with the free Harish-Chandra bimodule $\mathrm{U}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ associated to the $k$-th symmetric power. By restricting the action on the highest-weight vector for all $k$, one obtains a map of vector spaces
@@
where the target is the quantum Hamiltonian reduction of the semi-direct product with $\mathfrak{gl}_{N}$ acting on the symmetric algebra $S^{\bullet}\mathbf{C}^{N}$ in the standard way. In particular, it is an algebra as well. If we denote by $\mathfrak{m}_{N}$ the mirabolic subalgebra of the subgroup of $\mathrm{GL}_{N}$ preserving the dual of the last basis vector in $\mathbf{C}^{N}$ (see Section 6.3 for precise definition), then it turns out that this map restricts to a similarly defined quantum Hamiltonian reduction
$^{N}\!\mathcal{S}^{\mathrm{rat}}\rightarrow\mathrm{U}(\mathfrak{m}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}.$ (1.3)
The following is Theorems 6.14 and 6.9; here $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})^{\geq 0}$ is the non-negative part of the shifted Yangian. In fact, the explicit formulas for the second homomorphism are close to another presentation of the quantization of Zastava spaces for $\mathfrak{sl}_{2}$ from [15, Section 3.22].
@@
The map $^{N}\!\mathcal{S}^{\mathrm{rat}}\xrightarrow{\sim}\mathrm{U}(\mathfrak{m}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ is an algebra isomorphism. It can be extended to a surjective algebra homomorphism $\mathrm{Y}_{-2N}(\mathfrak{sl}_{2})^{\geq 0}\rightarrow\mathrm{U}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$.
Unfortunately, we do not know if and how this result can be related to the aforementioned stable bundles on an elliptic curve. We expect (or, rather, hope) that it could viewed as a quantum rational version of a certain algebra of Hecke modifications of semistable bundles of degree 0, see Remark 6.15. In general, it would be interesting to interpret this result at least on the classical level in terms of elliptic analog of the Kostant slice [18].
Besides degenerate affine Hecke algebras, the other main tool to prove the theorem is the Kirillov projector $P_{\psi}$ introduced in [27]. This is an element in a certain completion of $\mathrm{U}(\mathfrak{m}_{N})$ that acts on Whittaker modules and projects onto the space of Whittaker vectors (that is, on which $\mathfrak{n}_{+}$ acts with character $\psi$); as such, it is an analog of the classical extremal projector [1] in the case of trivial character. For instance, the Kirillov projector gives a canonical isomorphism $V\xrightarrow{\sim}\mathrm{U}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+}$ of any $\mathrm{GL}_{N}$-representation $V$ with the mirabolic version of the Kostant–Whittaker reduction. It turns out that this isomorphism has rather peculiar properties with respect to the degenerate affine Hecke algebra action mentioned above, see Section 6.3. For example, under the isomorphism (1.3), the standard monomial basis of $S^{\bullet}\mathbf{C}^{N}$ is mapped to a rational version of the Hall–Littlewood polynomials, see Example 6.12.
We extend this homomorphism to the Toda lattice. Naturally, $\mathrm{GL}_{N}$ is embedded into the matrix space; denote by $X=(x_{ji})$ the (transposed) matrix with $x_{ij}\in\mathcal{O}(\mathrm{GL}_{N})$ the $(i,j)$-matrix entry function. Its inverse $X^{-1}$ is well-defined in $\mathcal{O}(\mathrm{GL}_{N})$. Also, denote by $E^{R}=\bigl(E_{ij}^{R}\bigr)$ the matrix composed of left-invariant vector fields associated to the standard generators of the Lie algebra $\mathfrak{gl}_{N}$. The following is Theorem 8.10. It would interesting to compare this explicit form with the approach of [14] via coproducts on both sides.
@@
$x_{+}(u)\mapsto-\sum_{i=0}^{\infty}\bigl[\bigl(E^{R}\bigr)^{i}X\bigr]_{1N}u^{-i-1},\qquad x_{-}(u)\mapsto\sum_{i=0}^{\infty}\bigl[X^{-1}\bigl(E^{R}\bigr)^{i}\bigr]_{1N}u^{-i-1},$
@@
The following is Theorem 8.8 and the second part of Theorem 8.10.